OpenStudy (anonymous):

The combined land area of the countries A and B is 157,973 square kilometers. Country A is larger by 773 square kilometers. Determine the land area of each country.

OpenStudy (anonymous):

So we know that A + B = 157, 973 If country A is larger we can say that it'll be represented as x If country B is smaller than country A, let's represent it as x - 773

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

SO. .... should I divide 157,973 in half then add 773 to one of them? and that equals A?

OpenStudy (anonymous):

You could. But the approach I think your teacher wants you to reach is A + B = 157973 Since A can be represented by " x " we sub that in and B can be represented by "x - 773" x + (x -773) = 157973 2x - 773 = 157973 2x = 157200 x = 78600 Therefore x - 773 = 78600 - 773= 77827

OpenStudy (anonymous):

I did 157,973 -773 = then divide by 2 +773 = Country A Then subracted A from 157,973 which = Country B. thanks for the help and understanding. :)

OpenStudy (anonymous):

No problem c: Glad I could help